# Tree-level 11D supergravity amplitudes from the pure spinor worldline

### v1

@@ 4. Tree-level N-point scattering amplitudes in 11D
where the sum runs over integers $k$ and corresponding multi-particle labels $P_{i}$ such that $P_{1}P_{2}\dots P_{k}=2\dots N-1$.
@@
$\eqalignno{&P_{a}\to-\partial_{a},\quad d_{\alpha}\to D_{\alpha},\quad\omega_{\alpha}\to-\partial_{\lambda^{\alpha}},&({4.2})}$
@@ 4.1. BRST and gauge invariance
$\eqalignno{{\cal A}_{5}&=\int_{\infty}^{\infty}d\tau_{2}\int_{-\infty}^{\infty}d\tau_{3}\left\langle U_{1}^{(3)}(\infty)U_{2}^{(0)}(\tau_{2})U_{3}^{(0)}(\tau_{3})U_{4}^{(1)}(0)U_{5}^{(3)}(-\infty)\right\rangle&\cr&+\int_{-\infty}^{\infty}d\tau\left\langle U_{1}^{(3)}(\infty)U_{2}^{(0)}(\tau)V_{34}^{(1)}(0)U_{5}^{(3)}(-\infty)\right\rangle+\int_{-\infty}^{\infty}d\tau\left\langle U_{1}^{(3)}(\infty)U_{3}^{(0)}(\tau)V_{24}^{(1)}(0)U_{5}^{(3)}(-\infty)\right\rangle\cr&+\int_{-\infty}^{\infty}d\tau\left\langle U_{1}^{(3)}(\infty)U_{23}^{(0)}(\tau)U_{4}^{(1)}(0)U_{5}^{(3)}(-\infty)\right\rangle+\left\langle U_{1}^{(3)}(\infty)V_{234}^{(1)}(0)U_{5}^{(3)}(-\infty)\right\rangle.&({4.3})}$
@@
$\eqalignno{-&\int_{\infty}^{\infty}d\tau_{2}\int_{-\infty}^{\infty}d\tau_{3}\left\langle U_{1}^{(3)}(\infty)\left(\partial_{\tau_{2}}U_{2}^{(1)}(\tau_{2})U_{3}^{(0)}(\tau_{3})+U_{2}^{(0)}(\tau_{2})\partial_{\tau_{3}}U_{3}^{(1)}(\tau_{3})\right)U_{4}^{(1)}(0)U_{5}^{(3)}(-\infty)\right\rangle\cr=&\int_{-\infty}^{\infty}d\tau\left\langle U_{1}^{(3)}(\infty)\left(U_{2}^{(1)}(\tau+\epsilon)U_{3}^{(0)}(\tau)-U_{2}^{(0)}(\tau)U_{3}^{(1)}(\tau-\epsilon)\right)V_{4}^{(1)}(0)U_{5}^{(3)}(-\infty)\right\rangle\cr+&\int_{-\infty}^{\infty}d\tau_{3}\left\langle U_{1}^{(3)}(\infty)\left(U_{2}^{(1)}(\epsilon)U_{4}^{(1)}(0)-U_{2}^{(1)}(-\epsilon)U_{4}^{(1)}(0)\right)U_{3}^{(0)}(\tau_{3})U_{5}^{(3)}(-\infty)\right\rangle&\cr+&\int_{-\infty}^{\infty}d\tau_{2}\left\langle U_{1}^{(3)}(\infty)U_{2}^{(0)}(\tau_{2})\left(U_{3}^{(1)}(\epsilon)U_{4}^{(1)}(0)-U_{3}^{(1)}(-\epsilon)U_{4}^{(1)}(0)\right)U_{5}^{(3)}(-\infty)\right\rangle,&({4.4})}$
where $\epsilon$ represents an arbitrary small parameter, and we dropped the terms where $U^{(1)}$ approaches $\pm\infty$. After appropriate normal ordering, the parentheses in the second line of eqn. (4.4) can be written as an operator insertion
$\eqalignno{&U_{2}^{(1)}(\tau+\epsilon)U_{3}^{(0)}(\tau)-U_{2}^{(0)}(\tau)U_{3}^{(1)}(\tau-\epsilon)&\cr=&U^{(1)}_{2}U^{(0)}_{3}(\tau)-U^{(0)}_{2}U^{(1)}_{3}(\tau)+{1\over 2}\Big{(}[U^{(0)}_{2},U^{(1)}_{3}]+[U^{(0)}_{3},U^{(1)}_{2}]\Big{)}(\tau)\cr=&U^{(1)}_{[2}U^{(0)}_{3]}(\tau)+{1\over 2}[U^{(0)}_{(2},U^{(1)}_{3)}](\tau).&({4.5})}$
@@
$\eqalignno{&\int_{-\infty}^{\infty}d\tau\left\langle U_{1}^{(3)}(\infty)\left(U^{(1)}_{[2}U^{(0)}_{3]}+{1\over 2}[U^{(0)}_{(2},U^{(1)}_{3)}]\right)(\tau)V_{4}^{(1)}(0)U_{5}^{(3)}(-\infty)\right\rangle\cr+&\int_{-\infty}^{\infty}d\tau_{3}\left\langle U_{1}^{(3)}(\infty)\left\{U_{2}^{(1)},U_{4}^{(1)}\right\}(0)U_{3}^{(0)}(\tau_{3})U_{5}^{(3)}(-\infty)\right\rangle\cr+&\int_{-\infty}^{\infty}d\tau_{2}\left\langle U_{1}^{(3)}(\infty)U_{2}^{(0)}(\tau_{2})\left\{U_{3}^{(1)},U_{4}^{(1)}\right\}(0)U_{5}^{(3)}(-\infty)\right\rangle.&({4.6})}$
@@
$\eqalignno{&-\int_{-\infty}^{\infty}d\tau\left\langle U_{1}^{(3)}(\infty)U_{2}^{(0)}(\tau)\left\{U_{3}^{(1)},U_{4}^{(1)}\right\}(0)U_{5}^{(3)}(-\infty)\right\rangle\cr&-\left\langle U_{1}^{(3)}(\infty)\left(U_{2}^{(1)}(\epsilon)V_{34}^{(1)}(0)-U_{2}^{(1)}(-\epsilon)V_{34}^{(1)}(0)\right)U_{5}^{(3)}(-\infty)\right\rangle\cr&-\int_{-\infty}^{\infty}d\tau\left\langle U_{1}^{(3)}(\infty)U_{3}^{(0)}(\tau)\left\{U_{2}^{(1)},U_{4}^{(1)}\right\}(0)U_{5}^{(3)}(-\infty)\right\rangle&\cr&-\left\langle U_{1}^{(3)}(\infty)\left(U_{3}^{(1)}(\epsilon)V_{24}^{(1)}(0)-U_{3}^{(1)}(-\epsilon)V_{24}^{(1)}(0)\right)U_{5}^{(3)}(-\infty)\right\rangle,&({4.7})}$
@@
$\eqalignno{&\int_{-\infty}^{\infty}d\tau\left\langle U_{1}^{(3)}(\infty)\left(U^{(1)}_{[2}U^{(0)}_{3]}+{1\over 2}[U^{(0)}_{(2},U^{(1)}_{3)}]\right)(\tau)V_{4}^{(1)}(0)U_{5}^{(3)}(-\infty)\right\rangle\cr-&\left\langle U_{1}^{(3)}(\infty)\left\{U_{2}^{(0)},V_{34}^{(1)}\right\}(0)U_{5}^{(3)}(-\infty)\right\rangle-\left\langle U_{1}^{(3)}(\infty)\left\{U_{3}^{(0)},V_{24}^{(1)}\right\}(0)U_{5}^{(3)}(-\infty)\right\rangle.&({4.8})}$
@@
$\eqalignno{-&\left\langle U_{1}^{(3)}(\infty)\left(V_{23}^{(1)}(\epsilon)U_{4}^{(1)}(0)-V_{23}^{(1)}(-\epsilon)U_{4}^{(1)}(0)\right)U_{5}^{(3)}(-\infty)\right\rangle\cr-&\int_{-\infty}^{\infty}d\tau\left\langle U_{1}^{(3)}(\infty)\left(U^{(1)}_{[2}U^{(0)}_{3]}+{1\over 2}[U^{(0)}_{(2},U^{(1)}_{3)}]\right)(\tau)U_{4}^{(0)}(0)U_{5}^{(3)}(-\infty)\right\rangle&\cr+&\left\langle U_{1}^{(3)}(\infty)\left(\{U_{2}^{(1)},V_{34}^{(1)}\}+\{U_{3}^{(1)},V_{24}^{(1)}\}+\{U_{4}^{(1)},V_{23}^{(1)}\}\right)(0)U_{5}^{(3)}(-\infty)\right\rangle.&({4.11})}$
@@
$\eqalignno{-&\left\langle U_{1}^{(3)}(\infty)\{U_{4}^{(1)},V_{23}^{(1)}\}(0)U_{5}^{(3)}(-\infty)\right\rangle\cr-&\int_{-\infty}^{\infty}d\tau\left\langle U_{1}^{(3)}(\infty)\left(U^{(1)}_{[2}U^{(0)}_{3]}+{1\over 2}[U^{(0)}_{(2},U^{(1)}_{3)}]\right)(\tau)U_{4}^{(0)}(0)U_{5}^{(3)}(-\infty)\right\rangle\cr+&\left\langle U_{1}^{(3)}(\infty)\left(\{U_{2}^{(1)},V_{34}^{(1)}\}+\{U_{3}^{(1)},V_{24}^{(1)}\}+\{U_{4}^{(1)},V_{23}^{(1)}\}\right)(0)U_{5}^{(3)}(-\infty)\right\rangle.&({4.12})}$
@@
$\eqalignno{&\delta U^{(3)}=\{Q,\Sigma\}\ ,\ \delta U_{4}^{(1)}=[Q,\Lambda_{4}]\ ,\ \delta U_{i}^{(0)}=\partial_{\tau}\Lambda_{i}+[Q,\Omega_{i}]\ ,\ i=2,3\cr&\delta U_{23}^{(0)}=[Q,\Omega_{23}]+\Lambda_{[2}U^{(0)}_{3]}+{1\over 2}[U^{(0)}_{(2},\Lambda_{3)}]+U^{(1)}_{[2}\Omega_{3]}+{1\over 2}\{\Omega_{(2},U^{(1)}_{3)}\},&\cr&\delta V_{i4}^{(1)}=[\Lambda_{4},U_{i}^{{(1)}}]+[\Lambda_{i},U_{4}^{(1)}]+[Q,\Omega_{i4}]\ ,\ i=2,3\cr&\delta V_{234}^{(1)}=[V_{23}^{{(1)}},\Lambda_{4}]+[\Lambda_{3},V_{24}^{{(1)}}]+[\Lambda_{2},V_{34}^{{(1)}}]+[U_{2}^{(1)},\Omega_{34}]+[U_{3}^{(1)},\Omega_{24}]+[Q,\Omega_{234}],&\cr&&({4.13})}$
for arbitrary functions $\Sigma,\Lambda_{4},\Lambda_{3},\Lambda_{2},\Omega_{2},\Omega_{23},\Omega_{24},\Omega_{34},\Omega_{234}$. The invariance of (4.3) under the gauge symmetry associated to the parameter $\Sigma$ immediately follows from the same argument as that used in the proof of BRST closure below eqn. (4.3).
@@
$\eqalignno{\delta_{\Lambda_{4}}{\cal A}_{5}=&\int_{-\infty}^{\infty}d\tau\left\langle U_{1}^{(3)}\left(U^{(1)}_{[2}U^{(0)}_{3]}+{1\over 2}[U^{(0)}_{(2},U^{(1)}_{3)}]\right)(\tau)\Lambda_{4}(0)U_{5}^{(3)}\right\rangle\cr+&\int_{-\infty}^{\infty}d\tau\left\langle U_{1}^{(3)}\left(U^{(0)}_{2}(\tau)[U^{(1)}_{3},\Lambda_{4}](0)+U^{(0)}_{3}(\tau)[U^{(1)}_{2},\Lambda_{4}](0)\right)U_{5}^{(3)}\right\rangle\cr-&\int_{-\infty}^{\infty}d\tau\left\langle U_{1}^{(3)}U_{2}^{(0)}(\tau)[U_{3}^{(1)},\Lambda_{4}](0)U_{5}^{(3)}\right\rangle-\int_{-\infty}^{\infty}d\tau\left\langle U_{1}^{(3)}U_{3}^{(0)}(\tau)[U_{2}^{(1)},\Lambda_{4}](0)U_{5}^{(3)}\right\rangle\cr-&\int_{-\infty}^{\infty}d\tau\left\langle U_{1}^{(3)}\left(U^{(1)}_{[2}U^{(0)}_{3]}+{1\over 2}[U^{(0)}_{(2},U^{(1)}_{3)}]\right)(\tau)\Lambda_{4}(0)U_{5}^{(3)}\right\rangle-\left\langle U_{1}^{(3)}[V^{(1)}_{23},\Lambda_{4}](0)U_{5}^{(3)}\right\rangle\cr&+\left\langle U_{1}^{(3)}[V_{23}^{(1)},\Lambda_{4}](0)U_{5}^{(3)}\right\rangle=0.&({4.14})}$
@@
$\eqalignno{\delta_{\Lambda_{i}}{\cal A}_{5}&=-\int_{-\infty}^{\infty}d\tau\left\langle U_{1}^{(3)}\left(\Lambda_{[2}U^{(0)}_{3]}+{1\over 2}[U^{(0)}_{(2},\Lambda_{3)}]\right)(\tau)U^{(1)}_{4}(0)U_{5}^{(3)}\right\rangle&\cr&-\int_{-\infty}^{\infty}d\tau\left\langle U_{1}^{(3)}\left(U^{(0)}_{2}(\tau)[\Lambda_{3},U^{(1)}_{4}](0)+U^{(0)}_{3}(\tau)[\Lambda_{2},U^{(1)}_{4}](0)\right)U_{5}^{(3)}\right\rangle\cr&-\left\langle U_{1}^{(3)}[\Lambda_{2},V_{34}^{(1)}](0)U_{5}^{(3)}\right\rangle-\left\langle U_{1}^{(3)}[\Lambda_{3},V_{24}^{(1)}](0)U_{5}^{(3)}\right\rangle\cr&+\int_{-\infty}^{\infty}d\tau\left\langle U_{1}^{(3)}U_{2}^{(0)}(\tau)[\Lambda_{3},U^{(1)}_{4}](0)U_{5}^{(3)}\right\rangle+\int_{-\infty}^{\infty}d\tau\left\langle U_{1}^{(3)}U_{3}^{(0)}(\tau)[\Lambda_{2},U_{4}^{(1)}](0)U_{5}^{(3)}\right\rangle\cr&+\int_{-\infty}^{\infty}d\tau\left\langle U_{1}^{(3)}\left(\Lambda_{[2}U^{(0)}_{3]}+{1\over 2}[U^{(0)}_{(2},\Lambda_{3)}]\right)(\tau)U_{4}^{(1)}(0)U_{5}^{(3)}\right\rangle\cr&+\left\langle U_{1}^{(3)}\left([\Lambda_{2},V_{34}^{(1)}]+[\Lambda_{3},V_{24}^{(1)}]\right)(0)U_{5}^{(3)}\right\rangle=0.&({4.15})}$
@@
$\eqalignno{\delta_{\Omega_{i}}{\cal A}_{5}=&\int_{-\infty}^{\infty}d\tau\left\langle U_{1}^{(3)}\left(U^{(1)}_{[2}\Omega_{3]}+{1\over 2}\{\Omega_{(2},U^{(1)}_{3)}\}\right)(\tau)U^{(1)}_{4}(0)U_{5}^{(3)}\right\rangle&\cr+&\int_{-\infty}^{\infty}d\tau\left\langle U_{1}^{(3)}\left(\Omega_{2}(\tau)\{U^{(1)}_{3},U^{(1)}_{4}\}(0)+\Omega_{3}(\tau)\{U^{(1)}_{2},U^{(1)}_{4}\}(0)\right)U_{5}^{(3)}\right\rangle\cr-&\int_{-\infty}^{\infty}d\tau\left\langle U_{1}^{(3)}\Omega_{2}(\tau)\{U_{3}^{(1)},U^{(1)}_{4}\}(0)U_{5}^{(3)}\right\rangle-\int_{-\infty}^{\infty}d\tau\left\langle U_{1}^{(3)}\Omega_{3}^{(0)}(\tau)\{U_{2}^{(1)},U^{(1)}_{4}\}(0)U_{5}^{(3)}\right\rangle\cr-&\int_{-\infty}^{\infty}d\tau\left\langle U_{1}^{(3)}\left(U^{(1)}_{[2}\Omega_{3]}+{1\over 2}\{\Omega_{(2},U^{(1)}_{3)}\}\right)(\tau)U^{(4)}_{4}(0)U_{5}^{(3)}\right\rangle=0.&({4.16})}$
The first two lines of eqn. (4.16) arise from the variation of the first line in eqn. (4.3), after applying appropriate normal ordering. The third line results from the variation of the second line in eqn. (4.3), after IBP the BRST operator. The fourth line reflects the gauge transformation of the first term in the third line of eqn. (4.3), again after suitable normal ordering. Finally, the last line corresponds to the gauge transformation of the final term in eqn. (4.3).
@@
$\eqalignno{\delta_{\Omega_{i4}}{\cal A}_{5}=&-\left\langle U_{1}^{(3)}[U_{2}^{(1)},\Omega_{34}](0)U_{5}^{(3)}\right\rangle+\left\langle U_{1}^{(3)}[U_{3}^{(1)},\Omega_{24}](0)U_{5}^{(3)}\right\rangle\cr&+\left\langle U_{1}^{(3)}\left([U_{2}^{(1)},\Omega_{34}]+[U_{3}^{(1)},\Omega_{24}]\right)(0)U_{5}^{(3)}\right\rangle=0.&({4.17})}$

### v2

@@ 4. Tree-level N-point scattering amplitudes in 11D
where the sum runs over integers $k\in\{1,\dots,N-2\}$ and corresponding multi-particle labels $P_{i}$, such that $P_{1}\cup P_{2}\cup\dots\cup P_{k}=\{2,\dots,N-1\}$, and $P_{k}\cap\{N-1\}\neq\emptyset$.
@@
$\eqalignno{&P_{a}\to\partial_{a},\quad d_{\alpha}\to D_{\alpha},\quad\omega_{\alpha}\to\partial_{\lambda^{\alpha}},&({4.2})}$
@@ 4.1. BRST and gauge invariance
$\eqalignno{{\cal A}_{5}&=\int_{-\infty}^{\infty}d\tau_{2}\int_{-\infty}^{\infty}d\tau_{3}\left\langle U_{1}^{(3)}(\infty)U_{2}^{(0)}(\tau_{2})U_{3}^{(0)}(\tau_{3})U_{4}^{(1)}(0)U_{5}^{(3)}(-\infty)\right\rangle&\cr&+\int_{-\infty}^{\infty}d\tau\left\langle U_{1}^{(3)}(\infty)U_{2}^{(0)}(\tau)V_{34}^{(1)}(0)U_{5}^{(3)}(-\infty)\right\rangle+\int_{-\infty}^{\infty}d\tau\left\langle U_{1}^{(3)}(\infty)U_{3}^{(0)}(\tau)V_{24}^{(1)}(0)U_{5}^{(3)}(-\infty)\right\rangle\cr&+\int_{-\infty}^{\infty}d\tau\left\langle U_{1}^{(3)}(\infty)U_{23}^{(0)}(\tau)U_{4}^{(1)}(0)U_{5}^{(3)}(-\infty)\right\rangle+\left\langle U_{1}^{(3)}(\infty)V_{234}^{(1)}(0)U_{5}^{(3)}(-\infty)\right\rangle.&({4.3})}$
@@
$\eqalignno{-&\int_{-\infty}^{\infty}d\tau_{2}\int_{-\infty}^{\infty}d\tau_{3}\left\langle U_{1}^{(3)}(\infty)\left(\partial_{\tau_{2}}U_{2}^{(1)}(\tau_{2})U_{3}^{(0)}(\tau_{3})+U_{2}^{(0)}(\tau_{2})\partial_{\tau_{3}}U_{3}^{(1)}(\tau_{3})\right)U_{4}^{(1)}(0)U_{5}^{(3)}(-\infty)\right\rangle\cr=&\int_{-\infty}^{\infty}d\tau\left\langle U_{1}^{(3)}(\infty)\left(U_{2}^{(1)}(\tau+\epsilon)U_{3}^{(0)}(\tau)-U_{2}^{(1)}(\tau-\epsilon)U_{3}^{(0)}(\tau)\right)U_{4}^{(1)}(0)U_{5}^{(3)}(-\infty)\right\rangle\cr+&\int_{-\infty}^{\infty}d\tau\left\langle U_{1}^{(3)}(\infty)\left(U_{3}^{(1)}(\tau+\epsilon)U_{2}^{(0)}(\tau)-U_{3}^{(1)}(\tau-\epsilon)U_{2}^{(0)}(\tau)\right)U_{4}^{(1)}(0)U_{5}^{(3)}(-\infty)\right\rangle\cr+&\int_{-\infty}^{\infty}d\tau_{3}\left\langle U_{1}^{(3)}(\infty)\left(U_{2}^{(1)}(\epsilon)U_{4}^{(1)}(0)-U_{2}^{(1)}(-\epsilon)U_{4}^{(1)}(0)\right)U_{3}^{(0)}(\tau_{3})U_{5}^{(3)}(-\infty)\right\rangle&\cr+&\int_{-\infty}^{\infty}d\tau_{2}\left\langle U_{1}^{(3)}(\infty)U_{2}^{(0)}(\tau_{2})\left(U_{3}^{(1)}(\epsilon)U_{4}^{(1)}(0)-U_{3}^{(1)}(-\epsilon)U_{4}^{(1)}(0)\right)U_{5}^{(3)}(-\infty)\right\rangle,&({4.4})}$
where $\epsilon$ represents an arbitrary small parameter, and we dropped the terms where $U^{(1)}$ approaches $\pm\infty$. After appropriate normal ordering, the parentheses in the second and third lines of eqn. (4.4) can be written as operator insertions
$\eqalignno{&U_{2}^{(1)}(\tau+\epsilon)U_{3}^{(0)}(\tau)-U_{2}^{(1)}(\tau-\epsilon)U_{3}^{(0)}(\tau)=[U^{(1)}_{2},U^{(0)}_{3}](\tau),&({4.5})\cr&U_{3}^{(1)}(\tau+\epsilon)U_{2}^{(0)}(\tau)-U_{3}^{(1)}(\tau-\epsilon)U_{2}^{(0)}(\tau)=[U^{(1)}_{3},U^{(0)}_{2}](\tau).&({4.6})}$
@@
$\eqalignno{&\int_{-\infty}^{\infty}d\tau\left\langle U_{1}^{(3)}(\infty)\left([U^{(1)}_{2},U^{(0)}_{3}]+[U^{(1)}_{3},U^{(0)}_{2}]\right)(\tau)U_{4}^{(1)}(0)U_{5}^{(3)}(-\infty)\right\rangle\cr+&\int_{-\infty}^{\infty}d\tau_{3}\left\langle U_{1}^{(3)}(\infty)\left\{U_{2}^{(1)},U_{4}^{(1)}\right\}(0)U_{3}^{(0)}(\tau_{3})U_{5}^{(3)}(-\infty)\right\rangle\cr+&\int_{-\infty}^{\infty}d\tau_{2}\left\langle U_{1}^{(3)}(\infty)U_{2}^{(0)}(\tau_{2})\left\{U_{3}^{(1)},U_{4}^{(1)}\right\}(0)U_{5}^{(3)}(-\infty)\right\rangle.&({4.7})}$
@@
$\eqalignno{&-\int_{-\infty}^{\infty}d\tau\left\langle U_{1}^{(3)}(\infty)U_{2}^{(0)}(\tau)\left\{U_{3}^{(1)},U_{4}^{(1)}\right\}(0)U_{5}^{(3)}(-\infty)\right\rangle\cr&-\left\langle U_{1}^{(3)}(\infty)\left(U_{2}^{(1)}(\epsilon)V_{34}^{(1)}(0)-U_{2}^{(1)}(-\epsilon)V_{34}^{(1)}(0)\right)U_{5}^{(3)}(-\infty)\right\rangle\cr&-\int_{-\infty}^{\infty}d\tau\left\langle U_{1}^{(3)}(\infty)U_{3}^{(0)}(\tau)\left\{U_{2}^{(1)},U_{4}^{(1)}\right\}(0)U_{5}^{(3)}(-\infty)\right\rangle&\cr&-\left\langle U_{1}^{(3)}(\infty)\left(U_{3}^{(1)}(\epsilon)V_{24}^{(1)}(0)-U_{3}^{(1)}(-\epsilon)V_{24}^{(1)}(0)\right)U_{5}^{(3)}(-\infty)\right\rangle,&({4.8})}$
@@
$\eqalignno{&\int_{-\infty}^{\infty}d\tau\left\langle U_{1}^{(3)}(\infty)\left([U^{(1)}_{2},U^{(0)}_{3}]+[U^{(1)}_{3},U^{(0)}_{2}]\right)(\tau)U_{4}^{(1)}(0)U_{5}^{(3)}(-\infty)\right\rangle\cr-&\left\langle U_{1}^{(3)}(\infty)\left\{U_{2}^{(1)},V_{34}^{(1)}\right\}(0)U_{5}^{(3)}(-\infty)\right\rangle-\left\langle U_{1}^{(3)}(\infty)\left\{U_{3}^{(1)},V_{24}^{(1)}\right\}(0)U_{5}^{(3)}(-\infty)\right\rangle.&({4.9})}$
@@
$\eqalignno{-&\left\langle U_{1}^{(3)}(\infty)\left(V_{23}^{(1)}(\epsilon)U_{4}^{(1)}(0)-V_{23}^{(1)}(-\epsilon)U_{4}^{(1)}(0)\right)U_{5}^{(3)}(-\infty)\right\rangle\cr-&\int_{-\infty}^{\infty}d\tau\left\langle U_{1}^{(3)}(\infty)\left([U^{(0)}_{2},U^{(1)}_{3}]+[U^{(0)}_{3},U^{(1)}_{2}]\right)(\tau)U_{4}^{(1)}(0)U_{5}^{(3)}(-\infty)\right\rangle&\cr+&\left\langle U_{1}^{(3)}(\infty)\left(\{U_{2}^{(1)},V_{34}^{(1)}\}+\{U_{3}^{(1)},V_{24}^{(1)}\}+\{U_{4}^{(1)},V_{23}^{(1)}\}\right)(0)U_{5}^{(3)}(-\infty)\right\rangle.&({4.12})}$
@@
$\eqalignno{-&\left\langle U_{1}^{(3)}(\infty)\{U_{4}^{(1)},V_{23}^{(1)}\}(0)U_{5}^{(3)}(-\infty)\right\rangle\cr-&\int_{-\infty}^{\infty}d\tau\left\langle U_{1}^{(3)}(\infty)\left([U^{(0)}_{2},U^{(1)}_{3}]+[U^{(0)}_{3},U^{(1)}_{2}]\right)(\tau)U_{4}^{(1)}(0)U_{5}^{(3)}(-\infty)\right\rangle\cr+&\left\langle U_{1}^{(3)}(\infty)\left(\{U_{2}^{(1)},V_{34}^{(1)}\}+\{U_{3}^{(1)},V_{24}^{(1)}\}+\{U_{4}^{(1)},V_{23}^{(1)}\}\right)(0)U_{5}^{(3)}(-\infty)\right\rangle.&({4.13})}$
@@
$\eqalignno{&\delta U^{(3)}=\{Q,\Sigma\}\ ,\ \delta U_{4}^{(1)}=[Q,\Lambda_{4}]\ ,\ \delta U_{i}^{(0)}=\partial_{\tau}\Lambda_{i}+[Q,\Omega_{i}]\ ,\ i=2,3\cr&\delta U_{23}^{(0)}=\partial_{\tau}\Omega_{23}+2[U^{(0)}_{(2},\Lambda_{3)}]+2\{\Omega_{(2},U^{(1)}_{3)}\}+[Q,\Sigma_{23}],&\cr&\delta V_{i4}^{(1)}=[\Lambda_{4},U_{i}^{{(1)}}]+[\Lambda_{i},U_{4}^{(1)}]+[Q,\Omega_{i4}]\ ,\ i=2,3\cr&\delta V_{234}^{(1)}=[V_{23}^{{(1)}},\Lambda_{4}]+[\Lambda_{3},V_{24}^{{(1)}}]+[\Lambda_{2},V_{34}^{{(1)}}]+[U_{2}^{(1)},\Omega_{34}]+[U_{3}^{(1)},\Omega_{24}]+[U_{4}^{(1)},\Omega_{23}]+[Q,\Omega_{234}],&\cr&&({4.14})}$
for arbitrary functions $\Sigma,\Sigma_{23},\Lambda_{4},\Lambda_{3},\Lambda_{2},\Omega_{2},\Omega_{23},\Omega_{24},\Omega_{34},\Omega_{234}$. The invariance of (4.3) under the gauge symmetry associated to the parameter $\Sigma$ immediately follows from the same argument as that used in the proof of BRST closure below eqn. (4.3).
@@
$\eqalignno{\delta_{\Lambda_{4}}{\cal A}_{5}=&\int_{-\infty}^{\infty}d\tau\left\langle U_{1}^{(3)}\left([U^{(0)}_{2},U^{(1)}_{3}]+[U^{(0)}_{3},U^{(1)}_{2}]\right)(\tau)\Lambda_{4}(0)U_{5}^{(3)}\right\rangle\cr+&\int_{-\infty}^{\infty}d\tau\left\langle U_{1}^{(3)}\left(U^{(0)}_{2}(\tau)[U^{(1)}_{3},\Lambda_{4}](0)+U^{(0)}_{3}(\tau)[U^{(1)}_{2},\Lambda_{4}](0)\right)U_{5}^{(3)}\right\rangle\cr-&\int_{-\infty}^{\infty}d\tau\left\langle U_{1}^{(3)}U_{2}^{(0)}(\tau)[U_{3}^{(1)},\Lambda_{4}](0)U_{5}^{(3)}\right\rangle-\int_{-\infty}^{\infty}d\tau\left\langle U_{1}^{(3)}U_{3}^{(0)}(\tau)[U_{2}^{(1)},\Lambda_{4}](0)U_{5}^{(3)}\right\rangle\cr-&\int_{-\infty}^{\infty}d\tau\left\langle U_{1}^{(3)}\left([U^{(0)}_{2},U^{(1)}_{3}]+[U^{(0)}_{3},U^{(1)}_{2}]\right)(\tau)\Lambda_{4}(0)U_{5}^{(3)}\right\rangle-\left\langle U_{1}^{(3)}[V^{(1)}_{23},\Lambda_{4}](0)U_{5}^{(3)}\right\rangle\cr&+\left\langle U_{1}^{(3)}[V_{23}^{(1)},\Lambda_{4}](0)U_{5}^{(3)}\right\rangle=0.&({4.15})}$
@@
$\eqalignno{\delta_{\Lambda_{i}}{\cal A}_{5}&=-2\int_{-\infty}^{\infty}d\tau\left\langle U_{1}^{(3)}[U^{(0)}_{(2},\Lambda_{3)}](\tau)U^{(1)}_{4}(0)U_{5}^{(3)}\right\rangle&\cr&-\int_{-\infty}^{\infty}d\tau\left\langle U_{1}^{(3)}\left(U^{(0)}_{2}(\tau)[\Lambda_{3},U^{(1)}_{4}](0)+U^{(0)}_{3}(\tau)[\Lambda_{2},U^{(1)}_{4}](0)\right)U_{5}^{(3)}\right\rangle\cr&-\left\langle U_{1}^{(3)}[\Lambda_{2},V_{34}^{(1)}](0)U_{5}^{(3)}\right\rangle-\left\langle U_{1}^{(3)}[\Lambda_{3},V_{24}^{(1)}](0)U_{5}^{(3)}\right\rangle\cr&+\int_{-\infty}^{\infty}d\tau\left\langle U_{1}^{(3)}U_{2}^{(0)}(\tau)[\Lambda_{3},U^{(1)}_{4}](0)U_{5}^{(3)}\right\rangle+\int_{-\infty}^{\infty}d\tau\left\langle U_{1}^{(3)}U_{3}^{(0)}(\tau)[\Lambda_{2},U_{4}^{(1)}](0)U_{5}^{(3)}\right\rangle\cr&+2\int_{-\infty}^{\infty}d\tau\left\langle U_{1}^{(3)}[U^{(0)}_{(2},\Lambda_{3)}](\tau)U_{4}^{(1)}(0)U_{5}^{(3)}\right\rangle\cr&+\left\langle U_{1}^{(3)}\left([\Lambda_{2},V_{34}^{(1)}]+[\Lambda_{3},V_{24}^{(1)}]\right)(0)U_{5}^{(3)}\right\rangle=0.&({4.16})}$
@@
$\eqalignno{\delta_{\Omega_{i}}{\cal A}_{5}=&2\int_{-\infty}^{\infty}d\tau\left\langle U_{1}^{(3)}\{\Omega_{(2},U^{(1)}_{3)}\}(\tau)U^{(1)}_{4}(0)U_{5}^{(3)}\right\rangle&\cr+&\int_{-\infty}^{\infty}d\tau\left\langle U_{1}^{(3)}\left(\Omega_{2}(\tau)\{U^{(1)}_{3},U^{(1)}_{4}\}(0)+\Omega_{3}(\tau)\{U^{(1)}_{2},U^{(1)}_{4}\}(0)\right)U_{5}^{(3)}\right\rangle\cr-&\int_{-\infty}^{\infty}d\tau\left\langle U_{1}^{(3)}\Omega_{2}(\tau)\{U_{3}^{(1)},U^{(1)}_{4}\}(0)U_{5}^{(3)}\right\rangle-\int_{-\infty}^{\infty}d\tau\left\langle U_{1}^{(3)}\Omega_{3}^{(0)}(\tau)\{U_{2}^{(1)},U^{(1)}_{4}\}(0)U_{5}^{(3)}\right\rangle\cr-&2\int_{-\infty}^{\infty}d\tau\left\langle U_{1}^{(3)}\{\Omega_{(2},U^{(1)}_{3)}\}(\tau)U^{(4)}_{1}(0)U_{5}^{(3)}\right\rangle=0.&({4.17})}$
The first two lines of eqn. (4.17) arise from the variation of the first line in eqn. (4.3), after applying appropriate normal ordering. The third line results from the variation of the second line in eqn. (4.3), after IBP the BRST operator. Finally, the fourth line reflects the gauge transformation of the first term in the third line of eqn. (4.3), again after suitable normal ordering.
@@
$\eqalignno{\delta_{\Omega_{i4}}{\cal A}_{5}=&-\left\langle U_{1}^{(3)}[U_{2}^{(1)},\Omega_{34}](0)U_{5}^{(3)}\right\rangle+\left\langle U_{1}^{(3)}[U_{3}^{(1)},\Omega_{24}](0)U_{5}^{(3)}\right\rangle\cr&+\left\langle U_{1}^{(3)}\left([U_{2}^{(1)},\Omega_{34}]+[U_{3}^{(1)},\Omega_{24}]\right)(0)U_{5}^{(3)}\right\rangle=0.&({4.18})}$